Sol Summary Report – 12 April 2018

Sol 4

Summary Title: Whatever It Takes

Author: Victoria Varone

Mission Status:

Sol Activity Summary: It was quite a busy day for Crew 192 as we spent almost all of our time literally “battening down the hatches” as severe winds caused the HAB some physical wear that needed repair and periodic observation.

It began at ~0220 when the wind peeked. We had to tie down the center wooden cap for the structure because it kept slamming against the ceiling and being pulled up by the wind. It is now secure.

We had an approved short morning scouting EVA …The commander recalled the team when the front airlock door vibrated and broke free from its hasp and the door swung open so hard that it broke the banister and damaged the habitat. The commander decided to have the team emergent repair the HAB by reattaching the hasp and screwing the banister back in place. The external repairs to the HAB must be made at a later date because they were not emergent and beyond the ability of the crew.

We completed another engineering EVA (ALL HANDS EFFORT) to clean the trailer water tank in the RAM so that our next crew can maintain a large and safe supply of water. This involved stringing 550 cord to make a safety line which allowed all personnel to get to the HAB without issue in the mounting wind.

Look Ahead Plan: The strong winds are forecast to continue into tomorrow and should slow down after that. Once it does, we will resume our exploration EVAs, continually exploring new sites and familiarizing ourselves further with ones we’ve already seen.

Anomalies in Work: None

Weather: Chilly and extremely windy. Winds blowing at approximately 30 mph with gusts 46 mph.

Crew Physical Status: All crew are physically healthy and safe.

Sol Summary – April 11th

Crew 192 Sol Summary Report

Sol 3

Summary Title: Gone With the Wind

Author: Victoria Varone

Mission Status: Today was the first really hot day of our mission, a surprise in between the last few days of cool weather and wind. We continued to explore further out into the Martian terrain, exercising extra caution this time to remain hydrated and not push ourselves past our limits in hazardous conditions. The search for our missing science experiments goes on and we simultaneously improvise new science experiments to execute in their place.

Sol Activity Summary: Our day started early with our first EVA, where crew members Joe Dituri and Andreaa Radulescu went to explore another site to do a bit of hiking and geological investigation. When they returned we all had lunch together, which was followed by a brainstorming session for our next few EVAs.

Then we relaxed and watched The Martian, which was surprisingly similar to our own mission (missing crew member, same number of remaining crew members, affinity for duck tape), especially since not long after, a crazy wind storm kicked up.

Upon hearing a lot of commotion outside, we decided to have crew members Ashok Narayanamoorthi and Richard Blakeman head out on an engineering EVA prior to our scheduled evening EVA to check all around the hab. They found that a railing near the main door was broken because the front air lock door swung open and smashed railing. They made a couple of adjustments to secure the remaining piece as well as the hab door. Following that, the wind became so strong and started to present a safety hazard to the crew so the leader of the away team (Richard Blakeman) terminated the EVA.

Look Ahead Plan: We will continue to extend our range of Martian exploration as far as we possibly can, but some forecast hazardous weather has prompted us to change up some of our EVA plans. With strong winds predicted for tomorrow and Friday, we may choose to instead remain closer to the hab and work on our water tank cleanup project in the RAM and resume exploration over the weekend. Safety ALWAYS comes first.

Anomalies in Work: None

Weather: Hot and windy. The temperature reached approximately 83 degrees today and the wind storm was severe.

Crew Physical Status: All crew are physically healthy and safe.

EVA: See EVA Report.

Reports to be filed: Sol Summary Report, GreenHab Report, EVA Request, EVA Report Journalist Report, Commander’s Report, Photo Report, Science Report

Support Requested: See Commander’s Report.

Sol Summary – April 10th

Crew 192 Sol 2 Summary Report

Sol 2

Summary Title: A Beautiful Day for EVAs

Author: Victoria Varone

Mission Status: It was a beautiful day for EVAs over here for Crew 192 on Mars. We continue to clean and organize things around the hab and have started getting more comfortable with EVAs, which will allow us to continue the search for fascinating natural wonders as well as our missing science experiments and crew member. The mission is going well.

Sol Activity Summary: The day started with a morning EVA, where three of our crew checked out the Marble Ritual site and surrounding area. They discovered some interesting geological structures and were particularly fascinated by a view of snow-capped mountains from a dusty, sandy red desert. After exploring, the crew returned to the hab, examined one of the hab’s water tanks for future cleaning, and headed inside for lunch.

After a great afternoon session of debriefing, brainstorming, and improvising for additional science experiments, the remaining crew members headed out for another EVA, coming back with excited descriptions of their hike in Kissing Camels Ridge. They also had some insight into the length of time of EVAs, noting that the amount of hiking they did in their 2.5 hour EVA expended just enough of their energy without causing serious dehydration. They’ve advise the rest of us to limit high-impact EVAs to about two and a half hours for safety.

We’re now resting and planning for tomorrow’s activities and may end the night with a space-based movie.

Look Ahead Plan: The goal is continue stretching our EVAs a little further as we keep searching for our missing science experiments and crew member. Based on tonight’s hiking EVA, we also plan to pay special attention to crew health on EVAs, making sure everyone is healthy and comfortable and no one becomes dehydrated and ill. A little attention is also being put towards possible activities for our Yuri’s Night celebration.

Anomalies in Work: None

Weather: The weather was beautiful again today, the skies started out clear and the weather cool, which worked well for the first EVA. It became overcast from the afternoon on, which allowed our second EVA crew members to explore the sites with different skies overhead.

Crew Physical Status: All crew are physically healthy and safe.

EVA: See EVA Report.

Reports to be filed: Sol Summary Report, GreenHab Report, EVA Request, Journalist Report, Commander’s Report, Photo Report

Support Requested: None

Sol Summary – April 9th

Crew 192 Sol 1 Summary Report

Sol 1

Summary Title: PHEnOM Blue Crew Has Landed

Author: Victoria Varone

Mission Status: Ground Control to Major Tom, PHEnOM Blue Crew 192 has landed at the Mars Desert Research Station. Unfortunately we’ve “lost” a crew-member during reentry to Mars. Our science adviser was accidentally jettisoned along with many of our scientific experiments. We have renamed our habitat the Bowie Base in honor of the inspiration from the “star Man” himself David Bowie. Our team’s mission is to overcome this adversity and persevere in the name of science. We have good intelligence that some of the projects may have survived reentry so we will be performing EVAs to scout possible locations where the science may have re-entered. We intend to perform pilot studies and other scientific information gathering techniques as well.

Sol Activity Summary: We started our day by making sure we have an adequate water supply for our mission. Following that, we scouted the area around our habitat and took some photos for documentation. We also checked on our remaining science experiments like the locking mechanism of our observatory, which eventually opened after several unsuccessful attempts.

Next, we tested our ATVs and rovers to make sure they’re working properly for our EVAs and scouting missions. It was interesting getting to use them on the dusty, sandy Martian landscape.

After our test, we headed back inside the hab for our first crew dinner during our mission, and after that, two crew members conducted our first short-range EVA to test our spacesuits and communications devices. Upon their return, our GreenHab Officer went out to water the plants and our XO conducted a short engineering EVA around the hab for some repairs.

Look Ahead Plan: As we continue to search for our lost science experiments, we’ll stretch our EVAs further and further out into the Martian landscape each day. We’re also experimenting with various types of dehydrated “space food” to see what brings us the best combination of nutrition, convenience, and satisfaction. Some of the dehydrated items have proven to be more difficult to cook with than expected.

Anomalies in Work: None

Weather: The weather was beautiful all day today, clear skies with very few clouds. We’re getting a good feel for the desert climate, it’s dryness, and its variation in temperature. It seems to be somewhat chilly in the morning, warmer by midday (even a bit hot), and cools down nicely at night, sometimes even becoming very cold during the night itself. The dry air is also a change for some of the crew.

Crew Physical Status: All crew are physically healthy and safe.

EVA: Today two crewmembers conducted one short-range EVA that lasted approximately one hour. See additional EVA report.

Reports to be filed: Sol Summary Report, HSO Report, GreenHab Report, EVA Request, Journalist Report, Commander’s Report, Photo Report

Support Requested: We would very much appreciate it if mission control could send us forecast sunset times (actual, civil, and nautical) for each day of our mission, which will allow us some better planning for EVAs.

Sol Summary – April 8th

Crew 192 Sol 0 Summary Report

Sol 0

Summary Title: Where the Hell is Hanksville?

Author: Victoria Varone

Mission Status: Crew 192’s mission is shaping up to be a great one. Today was largely spent preparing for our simulation, but the crew still managed to complete a handful of important tasks during their preparation and organization, such as some filling of the hab’s water tanks, a health/safety brief, and some pre-planning for their simulation.

Sol Activity Summary: On our first full day at MDRS, we “flipped this hab!” We cleaned the hab top the bottom, including lots of mud that was tracked in after last night’s surprise rain. We also organized everything in hab, sorting all supplies in each area of the hab.

Look Ahead Plan: The crew will finish up their pre-sim duties tomorrow morning and start their simulation tomorrow afternoon. The goal is complete a solid, uninterrupted sim, during which we will complete science experiments and explore personal and professionals goals.

Anomalies in Work: There was a loose external part of the hab that became very loud in hitting an adjacent beam in the midst of strong winds. The crew patched up the part and the hab is running smoothly.

Weather: It was sunny and warm most of the day, but some heavy winds picked up from around 1500 to approximately 1900.

Crew Physical Status: All crew are physically healthy and safe.

EVA: There were no EVAs done today as the crew is not yet in sim.

Reports to be filed: Sol Summary Report, HSO Report, GreenHab Report, EVA Request

Support Requested: None

Sol Summary – April 07th

Crew 191 Sol Summary Report
07 Apr 2018

Sol 14

Summary Title:
Cleaning and taking the HAB over to following crew
Miho Tsukishiro (XO)

Mission Status:
finish

Sol Activity Summary:
7:20 Breakfast
7:45 Briefing
9:00 EVA #17 The additional EVA (just for TV broadcasting)
13:00 Welcoming crew192
13:30 Lunch
14:00 Cleaning & Transportation of Water from Hollow Mountain
18:00 Gathering reports
19:00 CapCom

Look Ahead Plan:
leaving MDRS

Anomalies in work:
N/A

Weather:
Cloudy. Mild winds.

Crew Physical Status:
Good.

EVA:
N/A

Reports to be file:
– SOL Summary Report
– Engineer Report
– GreenHab Report
– Photo Report with Description

Support Requested:
N/A

Sol Summary – April 06th

Crew 191 Sol Summary Report
06 Apr 2018

Sol 13

Summary Title:
The last EVA and Breaking SIM!
Miho Tsukishiro (XO)

Mission Status:
finish simulation

Sol Activity Summary:
7:00 Breakfast
7:30 Briefing
10:00 EVA #16 The last EVA (leader: Yusuke Murakami, Miho Tsukishiro, Fumiei Morisawa) Habcom: Kai Takeda
13:00 Breaking SIM
15:00 Transportation of Water from Hollow Mountain
18:00 Gathering reports
19:00 CapCom

Look Ahead Plan:
additional EVA for TV broadcasting in Japan

Anomalies in work:
N/A

Weather:
Cloudy. Mild winds.

Crew Physical Status:
Good.

EVA:
See EVA Report.

Reports to be file:
– SOL Summary Report
– Engineer Report
– GreenHab Report
– Photo Report with Description
– EVA Report
– EVA Request
– Final Report
– Food Inventory

Support Requested:
N/A

Sol Summary – April 05th

Crew 191 Sol Summary Report
05 Apr 2018

Sol 12

Summary Title:
Drone-shooting and the “Plants Day”
Miho Tsukishiro (XO)

Mission Status:
in progress

Sol Activity Summary:
7:00 Breakfast
7:30 Briefing
9:30 EVA #14 3D Survey of the HAB by a Drone Camera (leader: Wataru Okamoto, Yusuke Murakami) Habcom: Fumiei Morisawa
11:45 Lunch in GreenHab
13:30 EVA #15 Observing Plants and Lichens (leader: Kai Takeda, Venzha Christ, Makoto Kawamura) Habcom: Fumiei Morisawa
14:00 Cleaning Water Tank
18:00 Gathering reports
19:00 CapCom

Look Ahead Plan:
See EVA Request.

Anomalies in work:
N/A

Weather:
Cloudy. Mild winds.

Crew Physical Status:
Good.

EVA:
See EVA Report.

Reports to be file:
– SOL Summary Report
– Engineer Report
– GreenHab Report
– Photo Report with Description
– EVA Report x2
– EVA Request
– Journalist Report

Support Requested:
N/A

Sol Summary – April 4th

Hi CapCom,

this is today’s Sol Summary Report.

Best,
Miho XO crew191

Crew 191 Sol Summary Report
04 Apr 2018

Sol 11

Summary Title:

Miho Tsukishiro (XO)

Mission Status:
in progress

Sol Activity Summary:
7:00 Breakfast
7:30 Briefing
9:00 EVA #12 CEMs (leader: Venzha Christ, Fumiei Morisawa, Wataru Okamoto) Habcom: Miho Tsukishiro
13:00 Lunch
14:00 EVA #13 Cleaning Water Tank (leader: Yusuke Murakami, Fumiei Morisawa, Makoto Kawamura) Habcom: Miho Tsukishiro
18:00 Gathering reports
19:00 CapCom

Look Ahead Plan:
See EVA Request.

Anomalies in work:
N/A

Weather:
Cloudy. Mild winds.

Crew Physical Status:
Good.

EVA:
See EVA Report.

Reports to be file:
– SOL Summary Report
– Engineer Report
– GreenHab Report
– Photo Report with Description
– EVA Report x2
– EVA Request
– Journalist Report

Support Requested:
N/A

Sol Summary – April 03rd

Crew 191 Sol Summary Report
03 Apr 2018

Sol 10

Summary Title:
First activity in RAM
Miho Tsukishiro (XO)

Mission Status:
in progress

Sol Activity Summary:
7:00 Breakfast
7:30 Briefing
9:00 EVA #9 CEM’s (leader: Venzha Christ, Miho Tsukishiro, Kai Takeda) & EVA #11 cleaning the white tank in RAM (Yusuke Murakami, Makoto Kawamura) Habcom: Fumiei Morisawa
12:30 Lunch
13:30 personal work
18:00 Gathering reports
19:00 CapCom

Look Ahead Plan:
EVA #12: CMEs EVA (Venzha’s Project)

Anomalies in work:
N/A

Weather:
Sunny. Mild winds.

Crew Physical Status:
Good.

EVA:
See EVA Report.

Reports to be file:
– SOL Summary Report
– Engineer Report
– GreenHab Report
– Photo Report with Description
– EVA Report x2
– EVA Request
– GreenHab Report

Support Requested:
N/A